Initially, water is at v = 0.04331 m3/kg and T = 250°C. Then, it changes its state in an isochoric process to T = 360°C. Its final phase is…

a.

liquid

b.

liquid-vapor

c.

vapor or gas

d.

unknown

e.

other

Initially, water is at v = 0.0013023 m3/kg and T = 250°C. Then, water is heated in an isochoric process to a saturated state. Its final state is the saturated vapor.

True

False

Thermodynamics Problem
Answer #1

ANSWER :

Consider the attached T-v diagram of water

And using any standard water Temperature table to get following values :

AT T = 250 °C , v( sat. liquid) = 0.001251 m3/kg, v(sat. vapor) = 0.05013 m3/kg

AT T = 360 °C v( sat. liquid) = 0.001892 m3/kg, v(sat. vapor) = 0.006945 m3/kg

Part a.)

note that initially the water is within the T-v dome (T = 250 °C and  v(sat. vapor) = 0.05013 m3/kg) and closer to the saturated vapor line (i.e.. the right side of the diagram) ((see attached image of T-v diagram)

now since process is isochoric so , At final stage :

T = 360 °C and  v = 0.04331 m3/kg

But at T = 360 °C , v(sat. vapor) = 0.006945 m3/kg

so, clearly water has crossed the saturated vapor line and entered the superheated vapor region.

THUS OPTION C.) IS CORRECT , i.e vapor or gas

(NOTE THAT IT HAS NOT CROESSED THE P = 22MPa line , so it has not entered the critical region , this can be shown easily by calculating the final pressure at this state)

Part b.)

Note that

AT T = 250 °C , v( sat. liquid) = 0.001251 m3/kg, v(sat. vapor) = 0.05013 m3/kg

But given volume = 0.0013023 m3/kg , which is slightly more than  v( sat. liquid) = 0.001251 m3/kg.

Thus , initially it is just inside the T-v dome and closer to the saturated liquid line

So, isochoric heating will be represented by a vertical line till it touches the saturated liquid line .

So, the Final state will be saturated liquid

THUS , CORRECT ANSWER IS FALSE

22 MPa. (Trans- critical (Subcooled liquid region ) I Region ) Critical Point Part a) (final stage) o (Supertreated saturated